# ALTo: Adaptive-Length Tokenizer for Autoregressive Mask Generation

### v1

@@ Abstract
While humans effortlessly draw visual objects and shapes by adaptively allocating attention based on their complexity, existing multimodal large language models (MLLMs) remain constrained by rigid token representations. Bridging this gap, we propose ALTo, an adaptive length tokenizer for autoregressive mask generation. To achieve this, a novel token length predictor is designed, along with a length regularization term and a differentiable token chunking strategy. We further build ALToLLM that seamlessly integrates ALTo into MLLM. Preferences on the trade-offs between mask quality and efficiency is implemented by group relative policy optimization (GRPO). Experiments demonstrate that ALToLLM achieves state-of-the-art performance with adaptive token cost on popular segmentation benchmarks. Code and models are released at https://github.com/yayafengzi/ALToLLM.
@@ 1 Introduction
Token Representation Flexibility Autonomous Adaptivity
@@
  • We develop ALToLLM, which integrates ALTo into a multimodal large language model (MLLM), enabling adaptive mask token generation for object segmentation tasks. The number of generated tokens can vary from as few as 2 to as many as 32, with most cases around 16, allowing ALToLLM to balance quality and efficiency under different scenarios via GRPO.
@@ 2 Related work
Visual tokenzers play important roles in various visual tasks, such as image reconstruction van2017vqvae ; esser2020taming , visual compression yan2025elastictok , and visual generation ramesh2021zero ; tian2024var ; lu2023unifiedio ; wang2025himtok . VQVAE van2017vqvae ; razavi2019generating and VQGANesser2020taming are popular frameworks that encode images into discrete 2D tokens by vector quantization. BEiT bao2021beit exploits visual tokens in masked image modeling. To reduce the redundancy in 2D space, TiTok yu2024titok and SEED ge2023planting produce 1D sequence for image tokenization. Methods above typically use a rigid number of tokens to represent images, regardless of the complexity of the visual content. To have flexible tokenization, FlexTok roman2025flextok projects images into 1D variable-length token sequences. ElasticTok yan2025elastictok proposed an adaptive tokenizer for images and videos by dropping a random number of latter tokens during training. ALIT duggal2025adaptive discretizes images into flexible-length tokens by recurrent distillation until the reconstruction quality is okay or meet the maximum iterations. HiMTok wang2025himtok learns 1D hierarchical mask tokens to represent segmentation masks from coarse to fine. However, these methods cannot decide an adaptive number of tokens autonomously. Trials have been made by heuristic rules about image reconstruction quality duggal2025adaptive ; yan2025elastictok , which increases computational overhead and becomes impossible for image generation tasks. Our proposed ALTo is both flexible and adaptive, as illustrated in Tab.  1.
@@
MLLM-based image segmentation. Recent MLLM-based segmentation methods primarily follow three paradigms lan2024text4seg ; wang2025himtok ; tang2025ufo . MLLM-segmentation joint models such as LISA lai2024lisa , GSVA xia2024gsva , GLaMM rasheed2024glamm , PixelLM ren2024pixellm , and PSALM zhang2024psalm , create semantic-to-pixel connections through LLM hidden states and rely on additional segmentation modules. Text-based methods, including Text4Seg lan2024text4seg , LLaFS zhu2024llafs and VistaLLM pramanick2024jack , represent masks as text sequences (pixel classes or polygon vertices), suffering from heuristic and inaccurate mask representation. Interestingly, segmentation masks could also be viewed as images so that we can rethink image segmentation as a mask generation task lu2023unifiedio ; Lu2023unifiedio2 ; bar2022visual . HiMTok wang2025himtok applies the idea by utilizing a hierarchical mask tokenizer into LLMs. Going a step further, ALToLLM generates adaptive-length token sequences, which is efficient and effective.
@@
Reinforcement learning. Reinforcement learning (RL) has become increasingly important for enhancing vision-language models setlur2024rl ; ma2024coevolving ; qu2025latent ; li2024getting ; zhai2025enhancing . Approaches like direct preference optimization rafailov2023direct and proximal policy optimization john2017ppo face challenges with data efficiency and reward stability. Group relative policy optimization (GRPO) shao2024deepseekmath has emerged as a promising alternative through its groupwise reward mechanism. Recent applications demonstrate GRPO’s effectiveness across various vision-language tasks. Visual-RFT liu2025visualrft combines GRPO with verifiable rewards for efficient model adaptation. Vision-R1 huang2025visionr1 employs GRPO with progressive thinking suppression for complex reasoning. Seg-Zero liu2025segzero achieves zero-shot segmentation through pure RL. These works apply RL to text output, while we make it for preference optimization on mask token output.
@@ 3.1 Overview
The proposed adaptive length tokenizer (ALTo) represents object masks as token sequences whose lengths adapt to the complexity of objects autonomously. Simple objects (e.g., a sphere) may require few tokens, while intricate structures (e.g., complicated shapes and multiple objects) may use up to 32 tokens. Built on this, ALToLLM is introduced to perform instruction mask generation for referring image segmentation, as shown in Fig. 3. We design a multi-stage training recipe for ALTo and ALToLLM to learn flexible, adaptive, and effective mask representations and achieve strong segmentation performance, as shown in Fig. 2.
@@ 3.2 ALTo
The novel token length predictor (TLP) determines the optimal number of tokens for each mask. TLP leverages the CLS token feature $T_{\text{cls}}$, which encodes global image features, together with the 32 mask token features $T\in\mathbb{R}^{32\times d}$, to predict a proper token length. $T_{\text{cls}}$ is used as a query in an attention mechanism to evaluate the importance of each mask token. For each mask token $T_{i}$, a gated key is generated by SwiGLU as $k_{i}=(W_{\text{v}}T_{i})\odot\sigma(W_{\text{g}}T_{i})$. The probability for each token to be the stopping point is computed via scaled dot-product attention: $p=\text{softmax}(q_{\text{cls}}k^{T}/\sqrt{d})$. The predicted length is computed as the mathematical expectation $\hat{L}=\sum_{i=1}^{32}i\cdot p_{i}$.
Accordingly, the first $\hat{L}$ tokens are selected and sent to the MD, while the remaining tokens are zero-padded, represented as $H\bigodot T$, where $H$ is a binary mask defined as $H=\mathbb{I}[i\leq\hat{L}]$. However, such token chunking strategy is not differentiable, which prevents gradients from the mask de-tokenizer from flowing back to the TLP. To address this, we introduce a differentiable token chunking strategy by considering the stopping probability distribution $p$. The probability that the $i$-th token is used is given by the cumulative probability $P_{i}=\sum_{j\geq i}p_{j}$, which indicates that the stop position is later than this token and provides a soft version of token chunking. This inspires us to apply a straight-through estimator as $\hat{T}=(P-\hat{P}+H)\bigodot T$. In this formulation, the predicted mask is then given by $M_{\text{pred}}=\text{MD}(\hat{T},X_{\text{img}})$.
In stage 1.5, we use a reconstruction loss $\mathcal{L}_{\text{Mask}}=\text{MSE}(M_{\text{pred}},M_{\text{gt}})$ to optimize mask reconstruction, and a length regularization term $\mathcal{L}_{\text{Length}}=-\lambda\hat{L}$ to encourage shorter token sequences, balancing accuracy and efficiency while MT and MD are frozen. The final combined loss is $\mathcal{L}_{\text{ALA}}=\mathcal{L}_{\text{Mask}}+\mathcal{L}_{\text{Length}}$. Further details about the length supervision design are provided in the Appendix. A.
Figure 3: Architecture of the proposed ALToLLM.
@@ 3.3 ALToLLM
[…]
@@
where $R_{\text{valid}}$ is $1$ if the sample strictly follows the format in Eq. 1, and 00 otherwise. $R_{\text{accuracy}}$ is the intersection-over-union (IoU) score between the predicted and ground truth masks, with the predicted mask reconstructed by the MD using the adaptive mask tokens. $R_{\text{efficiency}}$ is a linear penalty $-\alpha L_{i}$ proportional to the token length $L_{i}\in\{1,\ldots,32\}$, scaled by the trade-off parameter $\alpha\geq 0$.
@@ 4.1 Experimental settings
Implementation details. ALTo processes input and reconstructs masks at $256\times 256$ resolution. During training and inference, the MLLM processes images at $448\times 448$, while the pixel encoder encodes image at $1024\times 1024$. In stage 1, MT and MD are initialized from TiTok-L-32 yu2024titok , and the pixel encoder is initialized from SAM-ViT-L kirillov2023sam . In stage 1.5, the feature dimension of TLP is set to 1024, consistent with MT. The length penalty coefficient is set to 0.0001, 0.001, 0.01, or 0.1, among which 0.01 is found to be optimal in subsequent experiments and is chosen for later stages. In stage 2, ALToLLM-8B is initialized from InternVL-2.5-8B chen2024far . Stage 3 trains the RL model based on the stage 2 checkpoint, with the length penalty set to 1e-2, 5e-3, 3e-3, 2e-3, 1e-3, or 1e-4, which are compared in later experiments. The KL penalty is set to 1e-3. We sample 12 group responses with a temperature of 1 and top-k of 10. Stages 1, 1.5, and 3 are trained on 8$\times$×A100 GPUs (80GB each), and stage 2 on 16$\times$×A100 GPUs. Training durations are: stage 1 for 2 days, stage 1.5 for 3 hours, stage 2 for 5 days, and stage 3 for 5 hours.
@@ 4.2 Comparative results
We evaluate ALToLLM-8B on the following tasks, considering two variants: (1) a fixed-length version (ALToLLM-8B (FL)) using 32 mask tokens, and (2) an adaptive-length version (ALToLLM-8B (AL)) that dynamically generates 1–32 mask tokens.
@@
4.3 Adaptive length preference adjustment via reinforcement learning
@@
4.4 Ablation study on ALTo
We demonstrate that the pixel encoder is necessary for the reconstruction of complex masks. As shown in Table 7, removing the pixel encoder leads to a significant drop in reconstruction gIoU on the Multi-Target-SA1B validation dataset, indicating that pixel-level features are crucial for capturing fine-grained mask details. We compare different penalty coefficients during stage 1.5 training and analyze their impact on adaptive-length prediction to select an appropriate value, as shown in Fig. 8. The results indicate that a coefficient of 0.01 achieves a good balance between IoU and a wide range of predicted lengths, with high standard deviation and entropy. Therefore, we adopt this value for stage 2 training.
@@ 5 Conclusions
Appendix
@@ Appendix A Details of TLP
where $\mathcal{L}_{\text{Length}}=-\lambda\hat{L}$ penalizes longer token sequences. The gradient of the loss is given by
@@
To address this limitation, we introduce a differentiable token chunking strategy. Specifically, we use the cumulative probability $P_{i}=\sum_{j\geq i}p_{j}$ to represent the probability that the $i$-th token is used, and $\hat{P}$ denotes the detached version of $P$. We then construct a soft token chunking as $\hat{T}=(P-\hat{P}+H)\bigodot T$. In this way, we have $\nabla\hat{T}=\nabla P\bigodot T$. The overall gradient is given by
@@
$-\frac{\partial\mathcal{L}_{\text{mask}}}{\partial\hat{T}_{k+1}}T_{k+1}<\lambda<-\frac{\partial\mathcal{L}_{\text{mask}}}{\partial\hat{T}_{k}}T_{k}$
@@ Appendix B Prompt design
We prepared different prompt templates for instruction tuning on adaptive-length and fixed-length segmentation. Tabs 9 and 9 are the templates for adaptive-length ALToLLM tuning, and Tabs 11 and 11 are the templates for fixed-length version.
@@
Appendix C Application examples
Fig. 9 illustrates examples of ALToLLM in practical applications. As demonstrated, ALToLLM effectively segments the target object referred to by the user by leveraging adaptive-length mask tokens. This approach not only ensures high segmentation accuracy but also optimizes computational efficiency by dynamically adjusting the number of tokens based on object complexity.

### v2

@@ Abstract
While humans effortlessly draw visual objects and shapes by adaptively allocating attention based on their complexity, existing multimodal large language models (MLLMs) remain constrained by rigid token representations. Bridging this gap, we propose ALTo, an adaptive-length tokenizer for autoregressive mask generation. To achieve this, a novel token length predictor is designed, along with a length regularization term and a differentiable token chunking strategy. We further build ALToLLM that seamlessly integrates ALTo into MLLM. Preferences on the trade-offs between mask quality and efficiency is implemented by group relative policy optimization (GRPO). Experiments demonstrate that ALToLLM achieves state-of-the-art performance with adaptive token cost on popular segmentation benchmarks. Code and models are released at https://github.com/yayafengzi/ALToLLM.
@@ 1 Introduction
Token Representation Flexibility Autonomous
Adaptivity
@@
  • We develop ALToLLM, which integrates ALTo into a multimodal large language model (MLLM), enabling adaptive mask token generation for object segmentation tasks. The number of generated tokens can vary from as few as 2 to as many as 32, with most cases around 17, allowing ALToLLM to balance quality and efficiency under different scenarios via GRPO.
@@ 2 Related work
Visual tokenizers play important roles in various visual tasks, such as image reconstruction van2017vqvae; esser2020taming, visual compression yan2025elastictok, and visual generation ramesh2021zero; tian2024var; lu2023unifiedio; wang2025himtok. VQVAE van2017vqvae; razavi2019generating and VQGANesser2020taming are popular frameworks that encode images into discrete 2D tokens by vector quantization. BEiT bao2021beit exploits visual tokens in masked image modeling. To reduce the redundancy in 2D space, TiTok yu2024titok and SEED ge2023planting produce 1D sequence for image tokenization. Methods above typically use a rigid number of tokens to represent images, regardless of the complexity of the visual content. To have flexible tokenization, FlexTok roman2025flextok projects images into 1D variable-length token sequences. ElasticTok yan2025elastictok proposed an adaptive tokenizer for images and videos by dropping a random number of the latter tokens during training. ALIT duggal2025adaptive discretizes the images into flexible-length tokens by recurrent distillation until the reconstruction quality is good or the maximum iterations are met. HiMTok wang2025himtok learns 1D hierarchical mask tokens to represent coarse to fine segmentation masks. However, these methods cannot decide an adaptive number of tokens autonomously. Trials have been made by heuristic rules about image reconstruction quality duggal2025adaptive; yan2025elastictok, which increases computational overhead and becomes impossible for image generation tasks. Our proposed ALTo is both flexible and adaptive, as illustrated in Table 1.
@@
MLLM-based image segmentation methods primarily follow three paradigms lan2024text4seg; wang2025himtok; tang2025ufo. MLLM-segmentation joint models such as LISA lai2024lisa, GSVA xia2024gsva, GLaMM rasheed2024glamm, PixelLM ren2024pixellm, and PSALM zhang2024psalm, create semantic-to-pixel connections through LLM hidden states and rely on additional segmentation modules. Text-based methods, including Text4Seg lan2024text4seg, LLaFS zhu2024llafs and VistaLLM pramanick2024jack, represent masks as text sequences (pixel classes or polygon vertices), suffering from heuristic and inaccurate mask representation. Interestingly, segmentation masks could also be viewed as images so that we can rethink image segmentation as a mask generation task lu2023unifiedio; Lu2023unifiedio2; bar2022visual. HiMTok wang2025himtok applies the idea by utilizing a hierarchical mask tokenizer into LLMs. Going a step further, ALToLLM generates adaptive-length token sequences, which is efficient and effective.
@@
Reinforcement learning (RL) has become increasingly important for enhancing vision-language models setlur2024rl; ma2024coevolving; qu2025latent; li2024getting; zhai2025enhancing. Approaches like direct preference optimization rafailov2023direct and proximal policy optimization john2017ppo face challenges with data efficiency and reward stability. Group relative policy optimization (GRPO) shao2024deepseekmath has emerged as a promising alternative through its groupwise reward mechanism. Recent applications demonstrate GRPO’s effectiveness across various vision-language tasks. Visual-RFT liu2025visualrft combines GRPO with verifiable rewards for efficient model adaptation. Vision-R1 huang2025visionr1 employs GRPO with progressive thinking suppression for complex reasoning. Seg-Zero liu2025segzero achieves zero-shot segmentation through pure RL. These works apply RL to text output, while we make it for preference optimization on mask token output.
@@ 3.1 Overview
Figure 2: Architecture of the proposed ALToLLM.
The proposed adaptive-length tokenizer (ALTo) represents object masks as token sequences whose lengths adapt to the complexity of objects autonomously. Simple objects (e.g., a sphere) may require few tokens, while intricate structures (e.g., complicated shapes and multiple objects) may use up to 32 tokens. Built on this, ALToLLM is introduced to perform instruction mask generation for referring image segmentation, as shown in Fig. 2. We design a multi-stage training recipe for ALTo and ALToLLM to learn flexible, adaptive, and effective mask representations and achieve strong segmentation performance, as shown in Fig. 3.
@@ 3.2 ALTo
The novel TLP determines the optimal number of tokens for each mask. TLP leverages the CLS token feature $T_{\text{cls}}$, which encodes global image features, together with the 32 mask token features $T\in\mathbb{R}^{32\times d}$, to predict a proper token length. $T_{\text{cls}}$ is used as a query in an attention mechanism to evaluate the importance of each mask token. For each mask token $T_{i}$, a gated key is generated by SwiGLU as $k_{i}=(W_{\text{v}}T_{i})\odot\sigma(W_{\text{g}}T_{i})$. The probability for each token to be the stopping point is computed via scaled dot-product attention: $p=\text{softmax}(q_{\text{cls}}k^{T}/\sqrt{d})$. The predicted length is computed as the mathematical expectation $\hat{L}=\sum_{i=1}^{32}i\cdot p_{i}$.
Accordingly, the first $\hat{L}$ tokens are selected and sent to the MD, while the remaining tokens are zero-padded, represented as $H\bigodot T$, where $H$ is a binary mask defined as $H=\mathbb{I}[i\leq\hat{L}$]. However, such token chunking strategy is not differentiable, which prevents gradients from the mask de-tokenizer from flowing back to the TLP. To address this, we introduce a differentiable token chunking strategy by considering the stopping probability distribution $p$. The probability that the $i$-th token is used is given by the cumulative probability $P_{i}=1-\sum_{j<i}p_{j}$, which indicates that the stop position is later than this token and provides a soft version of token chunking. This inspires us to apply a straight-through estimator as $\hat{T}=(P-P.\text{detach}()+H)\bigodot T$. In this formulation, the predicted mask is then given by $M_{\text{pred}}=\text{MD}(\hat{T},X_{\text{img}})$.
In stage 1.5, we use a reconstruction loss $\mathcal{L}_{\text{Mask}}=\text{MSE}(M_{\text{pred}},M_{\text{gt}})$ to optimize mask reconstruction, and a length regularization term $\mathcal{L}_{\text{Length}}=\lambda\hat{L}$ to encourage shorter token sequences, balancing accuracy and efficiency while MT and MD are frozen. The final combined loss is $\mathcal{L}_{\text{ALA}}=\mathcal{L}_{\text{Mask}}+\mathcal{L}_{\text{Length}}$. Further details about the length supervision design are provided in the Appendix. A.
@@ 3.3 ALToLLM
$\texttt{<ALTo\_Start>}\underbrace{\texttt{<TOK}_{1}\texttt{>}\cdots\texttt{<TOK}_{L_{i}}\texttt{>}}_{L_{i}\text{ tokens}}\texttt{<ALTo\_End>},\quad L_{i}\in\{1,\ldots,32\}$ (1)
@@
where $R_{\text{valid}}$ is $1$ if the sample strictly follows the format in Eq. 1, and 0 otherwise. $R_{\text{accuracy}}$ is the intersection-over-union (IoU) score between the predicted and ground truth masks, with the predicted mask reconstructed by the MD using the adaptive mask tokens; it vanishes to 0 if any responses do not conform to the correct format. $R_{\text{efficiency}}$ is a linear penalty $\alpha L_{i}$ proportional to the token length $L_{i}\in\{1,\ldots,32\}$, scaled by the trade-off parameter $\alpha\geq 0$.
@@ 4.1 Experimental settings
[…]
@@ 4.2 Comparative results
We evaluate ALToLLM-8B on the following tasks, considering two variants in SFT: (1) a fixed-length version (ALToLLM-8B (FL)) using 32 mask tokens, and (2) an adaptive-length version (ALToLLM-8B (AL)) that dynamically generates 1–32 mask tokens.
@@
4.3 Adaptive-length preference adjustment via reinforcement learning
@@
4.4 Ablation Studies
Ablation on ALTo.
We first verify the necessity of the pixel encoder for reconstructing complex masks. As shown in Table 7, removing the pixel encoder leads to a substantial drop in reconstruction gIoU on the Multi-Target-SA1B validation set, confirming that pixel-level visual features are essential for capturing fine-grained mask details. We also study the effect of different penalty coefficients during Stage 1.5 training on adaptive-length mask prediction, as illustrated in Fig. 8. The results show that a coefficient of 0.01 strikes an optimal balance: it maintains high mask quality (in terms of IoU) while enabling a diverse range of predicted token lengths, as evidenced by high standard deviation and entropy in output lengths. We therefore adopt this value for Stage 2 training.
Ablation on ALToLLM.
To better understand the key components driving ALToLLM’s performance gains, we conduct comprehensive ablation studies on gRefCOCO liu2023gres, with results summarized in Table 8. Our analysis reveals that ALToLLM’s improvements in mask quality primarily stem from two factors: (1) pretraining on Multi-Target-SA1B, a dataset containing scenes with multiple and structurally complex objects, and (2) the pixel encoder, which provides high-resolution visual cues that refine mask boundary reconstruction. Moreover, the adaptive-length setting in SFT not only improves token efficiency—reducing the average output length by approximately 50%—but also slightly enhances mask quality. This demonstrates that adaptive token generation effectively eliminates redundancy while preserving, and even improving, mask quality. Finally, GRPO contributes marginally to absolute performance metrics but helps the model learn an effective trade-off between mask quality and token efficiency.
Table 8: Ablation study for ALToLLM on gRefCOCO validation dataset.
Components Metrics
Multi-Target-SA1B Pixel Encoder AL(SFT) GRPO cIoU gIoU Avg. Length
70.4 72.1 32.0
72.8 75.6 32.0
74.8 77.6 32.0
75.4 78.0 17.5
75.5 78.1 15.7
@@ 5 Conclusions
While ALTo effectively handles most segmentation tasks with adaptive token lengths (1-32 tokens), two key directions merit further exploration. First, our approach requires multiple training stages, which increases the engineering complexity. A simpler training pipeline is desirable for future applications. Second, though our pipeline is designed to be modality-agnostic (treating mask tokenization as a special case of image tokenization), our experiments currently focus on mask validation. Future work will extend ALTo to RGB image tokenization and validate its effectiveness across more general vision tasks.
@@ Appendix A Details of TLP
where $\mathcal{L}_{\text{Length}}=\lambda\hat{L}$ penalizes longer token sequences. The gradient of the loss is given by
@@
To address this limitation, we introduce a differentiable token chunking strategy. Specifically, we use the cumulative probability $P_{i}=1-\sum_{j<i}p_{j}$ to represent the probability that the $i$-th token is used, and $\hat{P}$ denotes the detached version of $P$. We then construct a soft token chunking as $\hat{T}=(P-\hat{P}+H)\bigodot T$. In this way, we have $\nabla\hat{T}=\nabla P\bigodot T$. The overall gradient is given by
@@
$\frac{\partial\mathcal{L}_{\text{mask}}}{\partial\hat{T}_{k+1}}T_{k+1}<\lambda<-\frac{\partial\mathcal{L}_{\text{mask}}}{\partial\hat{T}_{k}}T_{k}$
@@ Appendix B Prompt design
We prepared different prompt templates for instruction tuning on adaptive-length and fixed-length segmentation. Tabs 10 and 10 are the templates for adaptive-length ALToLLM tuning, and Tabs 12 and 12 are the templates for fixed-length version.
@@
Appendix C Results on reasoning segmentation
We evaluate ALToLLM-8B on the ReasonSeg benchmark [lai2024lisa], which demands complex visual reasoning for accurate segmentation. As shown in Table 13, our method achieves consistently superior performance in both cIoU and gIoU, demonstrating enhanced reasoning capabilities over the baseline.
Table 13: Reasoning segmentation results on ReasonSeg validation dataset.
Methods cIoU gIoU
LISA-7B [lai2024lisa] 46.0 44.4
LISA-7B (ft) [lai2024lisa] 54.0 52.9
SAM4MLLM-8B [chen2024sam4mllm] 60.4 58.4
HiMTok [wang2025himtok] 67.0 60.7
ALToLLM-8B 67.3 62.8
Appendix D Results on region understanding
Following the setup of GLaMM [rasheed2024glamm], we assess Region-Level Captioning on the RefCOCOg dataset [mao2016generation]. Our method outperforms GLaMM in both METEOR and CIDEr metrics, demonstrating enhanced ability to generate semantically accurate and descriptive captions for localized image regions.
Table 14: Region-level captioning results on RefCOCOg.
Methods METEOR CIDEr
GLaMM 16.2 106.0
ALToLLM-8B 16.5 110.1
Appendix E Application examples
Fig. 9 and Fig. 10 illustrates examples of ALToLLM in practical applications. As demonstrated, ALToLLM effectively segments the target object referred to by the user by leveraging adaptive-length mask tokens.
Figure 9: An example of complex scenarios